Sol Summary – March 13th

Crew 293 Sol Summary Report 13-03-2024

Sol: 24

Summary Title: Mission objective: Save the station

Author’s name: Marie Delaroche, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

No Sol on Mars is uneventful, especially when one gets closer and closer to the end of their mission… This morning, the last exploration EVA supported by 3D mapping took place at Kissing Camel Ridge, involving the Crew Engineer, Crew Journalist and Crew Scientist. As the Journalist and Engineer were searching for checkpoints, suit n°5 stopped working. While the Crew tried troubleshooting on site, they could not find the source of the issue, and the EVA had to be continued out of sim for one of the EVA crew members. Once they were all safely back at the Hab, the Crew Engineer was able to figure out the source of the issue.

Meanwhile, after doing some more troubleshooting on the Echofinder tool, we have decided to continue running the experiment without the AI assistant, using only Augmented Reality. One session was performed this afternoon, the final one will take place on Sol 25.

With the help of Mission Support, we were also able to test our Anomaly Monitoring Interface, while simultaneously performing a meaningful fix outside the station. At 16:08, a programmed depressurization alarm resounded, prompting the crew to follow a realistic emergency EVA procedure. The objective was to repair the tunnel to the Science Dome; the designated crew was successful, the tunnel is now much safer to travel through. The interface test was also successful. We believe this tool could be a great addition for future simulations.

Look Ahead Plan:

The final exploration EVA for the Photogrammetry EVA is to take place on Sol 25. We aim to dismantle the atmospheric instruments on Sol 26, before breaking sim at 6 PM.

Anomalies in work: EVA Suits 7, 9

Weather: Partly cloudy; low 5°C, high 15°C

Crew Physical Status: Nominal

EVA: Exploration of Kissing Camel Ridge (Photogrammetry)

Reports to be filed: Sol Summary, Journalist’s Report, GreenHab Report, Operations Report, Photos (6-8 pics), EVA Reports, EVA Request, Astronomy Report

Support Requested:

We would like to thank Ben for his support on this afternoon’s EVA, and for giving us the opportunity to perform it.

The tools used for the emergency EVA have been placed in the RAM Airlock. We will cut the protruding ends of the zip ties in the Science Dome tunnel tomorrow.

Copyright © The Mars Society. All rights reserved. | Main Site